OpenStudy (anonymous):

Multiply: 2 the square root of 11 • 8 the square root of 22 I understand to multiply 2*8 together and then 11*22 but after that the simplifying is what I don't understand

OpenStudy (anonymous):

8 the square root of 11 10 the square root of 33 64 the square root of 22 176 the square root of 2 these are the answer choices

OpenStudy (anonymous):

Yes, you got the first part right. 2*8 and sqrt(11*22) . Which gives you 16sqrt(242)

OpenStudy (anonymous):

so then how do I simplify that??

OpenStudy (anonymous):

the sqrt(242) can be writen as sqrt(2*11*11)

OpenStudy (anonymous):

I know the answer wont be either of the first 2

OpenStudy (zzr0ck3r):

\[2\sqrt{11}*8*\sqrt{22}=16\sqrt{11}*\sqrt{2}\sqrt{11}=11*16*\sqrt{2}=176\sqrt{2}\]

OpenStudy (anonymous):

so would the answer be \[176\sqrt{2}\]

OpenStudy (anonymous):

Yes, thats it

OpenStudy (zzr0ck3r):

correct